OpenStudy (anonymous):

Medal(: Write the equation of of a line with the points (2,2), (0,-2) in slope intercept form.

OpenStudy (anonymous):

y = mx + b

hartnn (hartnn):

do u know how to get the slope given 2 points ?

OpenStudy (anonymous):

The slope is 2 right?

hartnn (hartnn):

correct! :) so, m= 2 slope intercept form : y= mx+c now u just need the "c"

hartnn (hartnn):

c is actually the y- intercept, of the form (0,c) so 'look' at 2 points, i am sure you will be able to figure out the value of c

OpenStudy (anonymous):

-2?

hartnn (hartnn):

-2 is absolutely correct! another way to get 'c' is to plug in any of the points, like if i use (2,2) 2 = 2(2) + c giving c= -2 so, the equation y=mx+c , for this Q, will be y= 2x-2 any doubts ?

OpenStudy (anonymous):

Nope! Thanks so much!

hartnn (hartnn):

welcome ^_^